OpenStudy (anonymous):

Solving exponential and logarithmic equations: 3log6 (x+1)=9 i know i'm supposed to divide both sides by 3/log6 to get log(x+1)=3log6 (i think) but I'm not sure what i need to do next.

OpenStudy (anonymous):

the base is 6

OpenStudy (mathmale):

3log6 (x+1)=9 becomes \[\log_{6}(x+1)=3 \]

OpenStudy (anonymous):

@mathmale what happened to the 9?

OpenStudy (mathmale):

Use the left and right sides of this equation as the exponents of the base 6:\[6^{\log_{6}(x+1) }=6^3\]

OpenStudy (mathmale):

Dividing the original equation (both sides) by 3 reduces that 9 to a 3.

OpenStudy (mathmale):

Use the inverse property of the log and expo functions to simplify the above equation.

OpenStudy (mathmale):

Please note that \[10^{\log_{10}x }=x\]

OpenStudy (mathmale):

and follow that approach to tackle your math problem.

OpenStudy (anonymous):

ok so they should be equal when i'm finished?

OpenStudy (mathmale):

Please focus on the left side of your equation. The base is 6. Write the base 6 and then apply the given log to the base 6 of (x+1). This whole expression boils down to what final result?

OpenStudy (mathmale):